A 14.0 L gas cylinder is filled with 8.20 moles of gas. The tank is stored at 29 ∘C . What is the pressure in the tank?
Answer #1

Volume of gas, V=14.0L, No.of moles, n = 8.20, temperature of system, T = 29°C = (29+273) K = 302 K

pressure (P) of gas can be determined using ideal gas equation, PV = nRT, R= universal gas constant = 0.082 L.atm.mol-1.K-1

P = nRT/V = (8.20 moles*0.082 L.atm.mol-1.K-1 * 302K) / 14.0 L = 14.5 atm